# Canonical Policy: Learning Canonical 3D Representation for Equivariant Policy

### v1

@@ VII Conclusions and Limitations
In this paper, we present canonical policy, a principled framework for 3D equivariant imitation learning that unifies point cloud observations through a canonical representation. Built upon a rigorous theory of 3D canonical mappings, our method enables end-to-end learning of spatially equivariant policies from demonstrations. By leveraging geometric consistency through canonicalization and the expressiveness of generative policy models, such as diffusion models, canonical policy improves generalization and data efficiency in imitation learning.
We validate our approach on 12 diverse simulated tasks and 4 real-world manipulation tasks, covering 16 evaluation settings involving distribution shifts in object appearance, shape, camera viewpoint, and robot platform. Canonical policy consistently outperforms existing image-based and point cloud baselines, demonstrating superior generalization capability and data efficiency.
Despite its strong empirical performance, canonical policy has several limitations. First, both the Point Cloud Aggregation Encoder Block and the SO(3)-equivariant networks used in our framework rely on neighborhood-based geometric processing, which can be computationally expensive and time-consuming. This overhead may limit scalability in resource-constrained applications.

### v2

@@ VII Conclusions and Limitations
In this paper, we present canonical policy, a principled framework for 3D equivariant imitation learning that unifies point cloud observations through a canonical representation. Built upon a rigorous theory of 3D canonical mappings, our method enables end-to-end learning of spatially equivariant policies from demonstrations. By leveraging geometric consistency through canonicalization and the expressiveness of generative policy models, canonical policy improves generalization and data efficiency in imitation learning.
We validate our approach on 12 diverse simulated tasks and 4 real-world manipulation tasks, covering 16 evaluation settings involving distribution shifts in object appearance, shape, camera viewpoint, and robot platform. Canonical policy consistently outperforms existing image and point cloud baselines, demonstrating superior generalization capability and data efficiency.
Despite its strong empirical performance, canonical policy has several limitations. First, both the Point Cloud Aggregation Encoder Block and the SO(3)-equivariant networks used in our framework rely on neighborhood processing, which can be computationally expensive and time-consuming. This overhead may limit scalability in resource-constrained applications.